# revtex4-2Repair the float

### v1

@@ IV.1.1 Coarse-Graining The ION
In the previous section, we obtained the binary undirected adjacency matrix $\mathbf{A}^{(\ell=0)}$ representing the interactions among the 6-digits sectors (0-nodes). Here, we describe the coarse-graining procedure producing the multi-scale unfolding of $\mathbf{A}^{(0)}$.
@@
Secondly, we set an edge among two $\ell-nodes$ if there was at least one link among their higher resolution members. More formally, the coarse-grained $\mathbf{A}^{(\ell)}$ is calculated by applying the Equation 4. We chose this one-step procedure to easily generate every level without passing into the intermediate scales. However, the MSM accept every arbitrary-steps scheme, e.g. if $\ell=3$ then $0\to 1\to 3$ is also possible. Note that the coarse-grained graphs become fully-connected from $\ell=4$. Therefore, the statistical modelling would be possible up to $\ell=3$.
IV.2 World Trade Web
As a second application, we considered the World Trade Web (WTW) from the Gleditsch dataset [23], which reports the international trade flows (imports and exports) among all the world countries. We selected the year 2000 (the most recent one) and removed the states that were not reported in the BACI-CEPII GeoDist [24] as we would use the geographical distances to coarse-grain the WTW. This results in $N_{0}=185$ 0-nodes. Although we analyzed the year 2000 the methodology can be applied also to other years.
The dataset provides two columns which display, for every pair of nodes $(i,j)$: the export $w_{ij}$ and import $w_{ji}$ among $i,j$ in USA dollars. However, by flipping $i\leftrightarrow j$, the export $\omega_{ji}$ and import $\omega_{ij}$ don’t coincide with the previous value. Therefore, we used only the redefined $w_{ij}\leftarrow\frac{w_{ij}+\omega_{ji}}{2}$ [25] as the amount of trade from $i\textnormal{ to }j$.
As done for the ION, we symmetrized the connections by mapping $w_{ij}\to w^{{}^{\prime}}_{ij}:=\frac{1}{2}\left(w_{ij}+w_{ji}\right)$, i.e. the average flow between the two directions. By renaming $w^{{}^{\prime}}_{ij}$ as $w_{ij}$, it follows that the weights are symmetric, i.e. $w_{ij}=w_{ji}$. Note that the WTW has a high reciprocity of links, i.e. the connections in the WTW are almost always corresponded [26]. Therefore, its undirected approximation is a legitimate starting point to study the system. Moreover, we binarized the import-export matrix to get an adjacency matrix among the states. Formally, we projected all the positive values of the weighted matrix $\textbf{W}:=\left\{w_{ij}\right\}_{i\in[1,N_{0}]}$ to one, i.e. $\mathbf{A}:=\Theta(\textbf{W})$ where the $\Theta(\cdot)$ is the Heaviside function.
IV.2.1 Coarse-Graining of WTW
To coarse-grain the empirical WTW, assumed at level 00, we used the geographical distances [24] to iteratively merge “close” nodes into block-nodes as in [17]. Technically, this is done by means of a single-linkage agglomerative clustering model which returns a dendrogram where the leaves are the 0-nodes, the branching points are the block-countries, and the height of each branching point represents the distance, obtained via the single-linkage, between two leaves following the corresponding branches. Therefore, to calculate the partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$ we cut the dendrogram at $18$ hierarchical heights $\ell\in[0,17]$, such that the number of block-countries, namely $i_{\ell+1}:=\Omega_{\ell}(i_{\ell})$, are $N_{\ell}:=N_{0}-10\cdot\ell$. Note that the coarse-grained graphs become fully-connected from $\ell=7$. Therefore, the statistical modelling would be possible up to $\ell=6$.
V Methodology
In this section, we will define the scores used to evaluate the models over the ION and the WTW.
V.1 Embedding Dimension
The choice of the best embedding dimension is still an active Research topic [27, 28, 29]. Here, we relied on the “Minimum Description Length” principle approximated by the Bayesian Information Criteria (BIC) [28]. In the Supp.Mat. Table 1, we reported the BIC scores over the dimensions $D=1,2,8,16$ and also, for completeness, the AIC ones [27]. The best dimension, according to BIC, depends on the resolution levels: for the ION, $D^{best}_{\ell\geq 1}=1$ whereas $D^{best}_{\ell=0}=2$. Since the WTW is a less complex network, the BIC selects everytime the lowest dimension, i.e. $D^{best}_{\ell\geq 0}=1$ (see Supp. Mat.). This result could be expected since having more nodes, as for the lower levels, it implies more heterogeneity and, therefore, a bigger embedding dimension to grasp the ION. However, since this is the first time the MSM has been proposed, we will show the scores for all the $D=1,2,8,16$ dimensions.
V.2 Renormalization of LPCA
Having fitted the $\vec{b}_{i_{0}},\vec{c}_{i_{0}}$ (at level 00), in order to model $A^{(\ell)}$, one may either proceed with the Equation 8 or by refitting LPCA at that level. However, both solutions have similar drawbacks as there is no renormalization rule enforcing the scale-invariance (cfr. Equation 15). In the following, we will analyze them in detail.

### v2

@@ IV.2 Coarse-Graining The ION
In the previous section, we constructed the binary undirected adjacency matrix $\mathbf{A}^{(\ell=0)}$ representing the interactions among the 6-digits sectors (0-nodes). Here, we describe the coarse-graining procedure to obtain the multi-scale unfolding of $\mathbf{A}^{(0)}$.
At first, we grouped together the 0-nodes that share the same first 66-$\ell$ digits, with $\ell\in[0,4]$. For instance, the sectors $111191,111199$ are merged in the same community $11119$ starting from $\ell=1$.
Secondly, we set an edge among two $\ell$-nodes if there existed at least one connection between any of their constituent 0-nodes (see Equation 4). Our procedure uses a one-step approach to easily generate every level directly from the $\mathbf{A}^{(0)}$, without passing through the intermediate scales. The latter procedure is also allowed by the MSM.
Note that from $\ell=4$ onward, the coarse-grained graphs become fully-connected, limiting a meaningful statistical modelling up to $\ell=3$. For technical details on the evolution of the number of nodes $N_{\ell},$ links $L_{\ell}$, and density $\rho_{\ell}:=\frac{2L_{\ell}}{N_{\ell}\left(N_{\ell}-1\right)}$, refer to Table 2.
IV.3 World Trade Web
As a second application, we analyzed the World Trade Web (WTW) tracked by the Gleditsch dataset [28], which reports bilateral trade flows (imports and exports) for every country. We focused on the year 2000 (the most recent one), and excluded the missing states in the BACI-CEPII GeoDist [29], whose distances are required for the coarse-graining procedure. After this filtering step, the graph has $N_{0}=185$ 0-nodes. Although we illustrate the method for the year 2000, the methodology applies to any other period of time.
For every pair $(i,j)$, the dataset reports the exported volume $w_{ij}$ and the imported $\omega_{ji}$ one (all in USA dollars). Since the reported export from $i$ to $j$ ($w_{ij}$) generally differ from the import from $j$ to $i$ ($\omega_{ij}$), we redefined $w_{ij}\leftarrow\frac{w_{ij}+\omega_{ji}}{2}$ [30] as the averaged amount of trade from $i\textnormal{ to }j$.
Successively, as for the ION, we symmetrized the transaction matrix by mapping every weight $w_{ij}$ to the average flow between the two directions, namely $w_{ij}\to w^{{}^{\prime}}_{ij}:=\frac{1}{2}\left(w_{ij}+w_{ji}\right)$, i.e. the average flow between the two directions. By renaming $w^{{}^{\prime}}_{ij}$ as $w_{ij}$, the resulting weights are symmetric, i.e. $w_{ij}=w_{ji}$. Given the high link reciprocity of the transaction reported in the Gleditsch dataset [31], this undirected representation is a sound approximation. Then, we binarized the import-export matrix to construct the adjacency matrix $\mathbf{A}:=\Theta(\textbf{W})$ where the $\Theta(\cdot)$ is the Heaviside function.
IV.4 Coarse-Graining of WTW
To coarse-grain the level 0, we used the inter-country geographical distances [29] to iteratively merge closer nodes via single-linkage agglomerative clustering as in [17]. This procedure returns a dendrogram whose leaves are the 0-nodes, whose branching points are the block-countries, and whose branch heights equal the single-linkage distances between sub-clusters. By cutting this dendrogram at 18 successive levels $\ell\in[0,17]$ we obtain partitions ${\Omega_{\ell}}{\ell\geq 0}$, such that the number of block-countries are approximately $N_{\ell}:=N_{0}-10\cdot\ell$. For $\ell\geq 7$ the coarse-grained graph become fully connected. Therefore, a meaningful statistical modelling is possible up to $\ell=6$. See Table 2 for the evolution of the number of nodes, links, and network density.
V Results and Discussions
Here, we present the results of the LPCA and MSM models applied to the ION datasets. The WTW results are reported in the Appendix.
V.1 Scale-Invariance Evidence and Multi-Scale Clustering Coefficient
Figure 2: This figure illustrates the numerical evaluation of the two sides of Equation 34 at level $\ell=2$ for the $LPCA$ and MSMMSM models. More precisely, the left-hand side (LHS) is represented on the y-axis while the right-hand side (RHS) on the x-axis.
(a) Level 0, 2 - Binary Clustering Coefficient
Figure 3: Cross comparison of LPCA-(8,8) and MSM-16 in predicting the clustering coefficient (CC) subsection C.2. The upper panel reports the expected clustering coefficient at level 0, while the lower panel depicts its corresponding values at level 2. The first column refers to the LPCA-(8,8), whereas the second to the MSM-16.
(b) Level 2 - MSM-16 - Summed, Fitted, Observed Network Measurements
Figure 4: The plots display the predicted network measurements for the ION dataset at level 2 according to LPCA-(8,8) (upper plots) and the MSM-16 (lower plot). In each of the upper panels, the x-axis represents the observed measurements, whereas the y-axis shows the expected ones. From left to right, the plots illustrates the degree (DEG), the average-nearest-neighbor degree (ANND), and the clustering coefficient ($C$). The scatter plot comparing $\mathbf{P}_{sum}^{(2)}$ with $\hat{\mathbf{P}}^{(2)}$ is reported in the inset. The lower panels show the behavior of the network measurements as the degrees increase. The observed values are colored in blue, while the one calculated using the fitted $\hat{\mathbf{P}}^{(2)}$ model in red or the summed $\mathbf{P}_{sum}^{(2)}$ model in orange. Additionally, the z-score of the predicted number of links is indicated in the legend of the upper-left plot.
Figure 2 illustrates the relationship among the summed probability and its coarse-grained counterpart (see Equation 34, subsection C.1). We selected the lowest embedding dimension $D_{B}=D_{C}=1$ and $D=1$ to highlight the fundamental differences among the models. As shown, only the MSM recovers the identity line for all the pairs, demonstrating its scale-invariant property. In contrast, LPCA consistently underestimates the coarse-grained probability, since the microscopic probabilities, maximizing the likelihood at $\ell=0$, get lower values than required for scale-invariance.
Subsequently, we increased the embedding dimensions ($D_{B}=D_{C}=8$ and $D=16$) and cross-compared their capability of predicting the multi-level clustering coefficient (see subsection C.2) in Figure 3a. Visually, LPCA-(8,8) appears to (slightly) outperform the MSM-16 at level 0, while the MSM excels for $\ell=2$. Quantitatively, at $\ell=0$, LPCA achieves better metrics ($AUC-ROC_{LPCA}\approx 0.92,AUC-PR_{LPCA}\approx 0.89$) than the MSM ($AUC-ROC_{MSM}\approx 0.91,AUC-PR_{MSM}\approx 0.88$). To merge with [32], we exploited the relative Froebenius error between the fitted probability $\hat{\mathbf{P}}^{(0)}$ and the adjacency matrix, defined as [32]
$\epsilon_{2}:=\frac{\left\|\hat{\mathbf{P}}^{(2)}-\mathbf{A}^{(2)}\right\|_{2}}{L_{2}}.$ (18)
where $\left\|\right\|_{2}$ refers to the 2-Euclidean norm. Also in this metric, LPCA exploits a better approximation than the MSM ($\epsilon_{LPCA}\approx 48\%<\epsilon_{MSM}\approx 51\%$).
Comparable results were also obtained for the other levels and measures, such that the DEG and ANND - for ION at $\ell=2$ (see Figure 4), but that are not shown here for brevity. Refer to the Supplementary Material for the WTW results.
V.2 Expected Values of the Network Measurements
In Figure 4, we illustrate key network properties at level $2$ (see subsection C.2) computed for the empirical network $\mathbf{A}^{(2)}$, as well as for its expected values under the summed model $\mathbf{P}_{sum}^{(2)}$ and the fitted model $\hat{\mathbf{P}}^{(2)}$ (see section III and III.4.2). For the LPCA, we selected $D_{B}=8,D_{C}=8$, and for the MSM, $D=16$.
The empirical quantities were computed as in subsection C.2. Focusing on the fitted model $\hat{\mathbf{P}}^{(\ell=2)}$, we sampled $N_{\mathcal{A}}=1000$ realizations to
compute the ensemble averages $\left\{\overline{k_{i_{2}}},\overline{annd_{i_{2}}},\overline{c_{i_{2}}}\right\}_{i_{2}\in[1,N_{2}]}$ as defined in Equation 73. Furthermore, we calculated the dispersion interval for each measurement setting $c=95\%$, and represented as error bars in the plots. The same procedure was applied the summed model $\mathbf{P}_{sum}^{(2)}$ for comparison.
In the upper panel, we compare the observed measurements (x-axis) with their expected values (y-axis) for the summed model (orange points) and the fitted model (red points). The identity line represents the perfect agreement between observed and predicted measures. The inset displays a scatter plot between $\mathbf{P}_{sum}^{(2)}$ and $\hat{\mathbf{P}}^{(2)}$, as discussed in subsection V.1. The $z_{L}$ value is defined as the z-score of links:
$z_{L_{\ell}}:=\frac{L_{\ell}-\langle L_{\ell}\rangle}{\sigma\left[L_{\ell}\right]}$ (19)
where $\sigma\left[L_{\ell}\right]:=\left[\sum_{i<j}p_{ij}\left(1-p_{ij}\right)\right]^{1/2}$ [33].
As shown in Figure 2, the LPCA does not accurately reproduce the empirical measurements. In contrast, the MSM provides a better approximation, with most empirical values falling within the dispersion intervals. Moreover, the inset in the upper-left plot reveals that $\mathbf{P}_{sum}^{(2)}$ closely approximates $\hat{\mathbf{P}}^{(2)}$ only for the MSM. This result, which was not theoretically enforced as for Equation 34, accounts for a consistent agreement of the MSM with observed values across scales.
The lower panel illustrates how network measurements evolve as a function of the degree. From left to right, we show the complementary cumulative distribution function (CCDF) of the degrees, the average nearest neighbors degree (ANND) and the clustering coefficient (CC) subsection C.2. The decreasing trend of the empirical CCDF suggests that the observed network contains more low-degree nodes than the hubs, and it is not scale-free, as the CCDF doesn’t follow a straight line in log-log scale. Additionally, by looking at the ANND, the ION is disassortative, since high-degree nodes tend to trade with loosely-interacting partners [34].
These plots highlight that the MSM-16 not only captures the CC (see Figure 3a), but also the lower-hops structural properties. As expected, the LPCA-(8,8) provides a good fit only at the specific scale at which it was fitted.
V.3 Reconstruction Accuracy and ROC-PR Curves
(a) Level 2 - LPCA-(8,8) - ROC and PR curves for Summed and Fitted models
(b) Level 2 - MSM-16 - ROC and PR curves for Summed and Fitted models
Figure 5: The graphs show the Confusion matrices, ROC and PR curves for the ION dataset at level 2 obtained by means of the LPCA-(8,8) (upper plots) and the MSM-16 (lower plot). On the left most side, one may display the two confusion matrices for the fitted $\hat{\mathbf{P}}^{(2)}$ (upper) and the summed $\mathbf{P}_{sum}^{(2)}$. The middle plot reports the Receiver-Operator Curve, while the right-most plot depicts the Precision Recall curve. As in Figure 4, the two curves are associated with the fitted (red) model and the summed (orange) one.
(a) All level Reconstruction Accuracy
(b) AUC-ROC and AUC-PR curves through levels
Figure 6: Left: Reconstruction Accuracy (y-axis) by model, level and network statistics. Right: Area Under the ROC and PR curves (y-axis) for the summed models as the scales increase (x-axis) or, equivalently, the numbers of nodes diminish.
In Figure 6a, we show the reconstruction accuracy (see Equation 77) across the available levels ($\ell=0,\cdots,3$) of the ION for the three key topological measures (DEG, ANND, CC). For this exercise, we report only the summed LPCA with dimensions $D_{B}=D_{C}=1$ and $D_{B}=D_{C}=8$, and the summed MSM with $D=1,2,8,16$. As anticipated, all the models are fitted at $\ell=0$, which explains why most trends are peaked at this resolution. For example, MSM-1 has a higher CC at level $1$ than at 0. Interestingly, LPCA struggles to produce dispersion intervals that include the observed quantities. Contrarily, the MSM preserve consistency across all levels -except for $\ell=2$ where the observed topology changed more than expected by the model. This fact is more noticeable for $D=1$, where the MSM overestimates all the measurements, since the summation of the 0-parameters lead to bigger values than requested parameters at level $2$. However, this is not inherently problematic, since the scale-invariance property enforced the internal consistency (see Equation 34) over exact recovery of the fitted parameters at every level. Also, the choice of a pathological partition would lead to worse results as discussed in [35] and subsubsection V.3.1.
Increasing the number of parameters ($D$) generally improves reconstruction accuracy. However, this comes at the price of “overfitting”, as indicated by the higher BIC scores reported in Supplemental Material [20]. A potential solution for this mismatch could be to introduce dyadic relationships $d_{ij}$ between nodes as done in [17]. Still, this remains beyond the scope of the present work.
Similar conclusions can be drawn from Figure 6b, which displays the Area-Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) and the Precision-Recall (PR) curves, defined as in subsection C.2. These plots further highlights a form of phase separation due to different natures of the two model. While LPCA-(8,8) outperforms all the other candidates at $\ell=0$, its performances degrades at coarser levels, as does that of LPCA-(1,1). Conversely, the MSM displays increasing AUC scores at coarser levels. These increasing trends are mainly because, at coarser levels, the empirical number of links decreases $L_{\ell}$ (see Table 2), while the summed probabilities $p_{IJ}$. Due to the probabilities’ growth, also the thresholded Predicted Positives ($\langle L\rangle(\epsilon)$) increase, but at a lower rate than the $TP(\epsilon)$. As a result, both the $TPR(\epsilon)$ and $PPV(\epsilon)$ increase at coarser levels, leading to higher AUC scores.
Taken together, the results indicates, by selecting a specific functional form, one can prioritize either for single-scale performances with the LPCA or generalize at multiple scales with the MSM.
V.3.1 Dependence on an arbitrary partition
The agreement of the summed model and the observed (coarser) graph depends critically on the chosen partition $\Omega$. Formally, what has being tested so far is the approximation
$a_{IJ}$ $\approx p_{IJ}\stackrel{{\scriptstyle!}}{{=}}1-\prod_{i_{0}\in I,j_{0}\in J}\left(1-p\left(x_{i_{0}},x_{j_{0}},w_{i_{0}}\right)\right)$ (20)
which depends on the partition $\Omega$ since the block-nodes are defined as $I:=\Omega(i),J:=\Omega(j)$ (see Equation 1 and Equation 34 for the notation). As a result, the quality of the above approximation is sensitive of the partition. In fact, one can intentionally construct a partition $\Omega^{\text{diff}}$ that maximizes the discrepancy between the model probability $p_{IJ}$ and the observed adjacency $a_{IJ}$. Formally, this corresponds to set $\Omega^{diff}$ as
$\Omega^{diff}:=\max_{\Omega}\sum_{I\leq J}\left(1-\prod_{i_{0}\in I,j_{0}\in J}(1-p_{i_{0}j_{0}})-a_{IJ}\right)^{2}$ (21)
This formulation highlights the fact that poorly chosen partitions can lead to misleading conclusions regarding model fit and predictive accuracy. Hence, the meaningful agreement between the summed model and the observed adjacency matrix, we have appreciated so far, was not expected a priori.
V.4 Expected Number of Triangles
(a) Level 0
(b) Level 2
Figure 7: Left: Triangle Density at level 0. The plot shows the evolution of the $\tau^{(\ell)}(c)$ with respect to the degree $c$ in blue solid dots. The other markers identify the expected $\langle\tau^{(\ell)}(c)\rangle$ by the LPCA-(1,1), LPCA-(8,8), MSM-1, MSM-2, MSM-8, MSM-16. Right: Triangle Density at level 2.
The authors of [32] demonstrate that through LPCA, it is possible to reproduce the exact triangle density (TriDens) using a lower embedding dimension than the number of nodes (cfr. [16]). Contrarily, here we illustrate the expected TriDens, defined in Equation 75 and Equation 76.
In Figure 7, the filled blue markers depict the observed TriDens, while the other markers correspond to different models, as described in the legend. Figure 7a reports the case for level $\ell=0$ where it is clear that even the smallest embedding dimension ($D=1$) provides a good approximation of TriDens. By design, as the cut-off degree $c$ increases, the difference among the TriDens values diminish, eventually converging at c=N1c=N-1. This is because, in a disassortative network, the subgraph composed of hubs is unlikely to contain triangles (see Figure 4). This finding does not contradict previous studies [16, 32], since our analysis focuses on the expected TriDens rather than the exact quantities.
By coarse-graining the network Figure 7b at $\ell=2$, the MSM models show a good agreement with the coarser TriDens. In particular, as shown in Figure 4, the estimates increase with increasing embedding dimension $D$. In contrast, LPCA tend to underestimate the metrics at coarser level, since it is restricted to the fitted scale.
VI Conclusions
The strength of the graphs-based models lies in their flexibility to capture various type of interactions through suitable definitions of nodes and edges. By grouping microscopic nodes into a single block-node, the resulting network is a coarse-grained version of the original graph. Repeating this procedure several times, it produces a multi-scale and hierarchical unfolding of the observed graph. We applied this procedure to both the ION and WTW datasets (see subsection IV.1), demonstrating how a single generative process can be represented at different resolutions.
Most of the existing node-embedding methods [36] are designed to find the optimal node representation at a single scale, often underestimating the importance of extrapolating the model across different scales. To illustrate this limitation, we compared LPCA -a single-scale method from the machine-learning domain- with our vector-based multi-scale model (MSM) (section III). The crucial difference lies in the fact that the MSM is inherently scale-invariant, similarly to the (hidden) generative process it aims to approximate. In this context, a block embedding is computed as the sum of the embeddings of its constituent nodes (Equation 14), allowing for a principled interpretation of the sum of the embeddings.
Noticeably, the embeddings are generally thought to live in a vector space that naturally comes with the sum and difference operations. However, despite their relevance, these two operations have received limited attention in the literature. Here, instead, we were able to provide a solid explanation for the sum of vectors (see Equation 14).
As exposed in section V, LPCA may perform at its best only at the scale it is fitted to, and struggles to generalize at higher levels. In contrast, the multi-scale model consistently captures the network structure to coarser resolutions, requiring only the knowledge of how microscopic nodes are organized into communities.
Taken together, these results highlight the value of adopting a multi-scale perspective, which offers a flexible and robust framework for understanding complex systems by allowing arbitrary adjustment of the scale at which interactions are examined.
VI.1 Future Perspectives
Although this work focus on the undirected MSM, the framework can be readily adapted to capture other phenomena [37]. Specifically, each node is represented by two embeddings ($\vec{x}_{i},\vec{y}_{i}$), which in this context are roughly seen as the capability of “proposing” (sender) or “accepting” (receiver) an economic relationship, respectively [32]. As in the present essay, the block outer (inner) vectors are obtained by summing the $\vec{x}$ ($\vec{y}$) embeddings of its constituent nodes, ensuring that the sum of vectors remains interpretable.
For the weighted part, further theoretical investigation is required to determine how weights can be integrated into the MSM framework.
VII DATA AND CODE AVAILABILITY
The dataset on sector transactions (ION) used in the paper is highly confidential and cannot be disclosed. However, the world trade is a long-standing public dataset, one may find at Gleditsch. Lastly, the code is freely available as the multi-scale-node-embeddings python package.
VIII ACKNOWLEDGMENTS
We thank ING Bank N.V. for their support and active collaboration. A special thanks to the whole DataScience team at ING Bank for their advice that helped shape this research. This work is supported by the project “Reconstruction, Resilience and Recovery of Socio-Economic Networks” RECON-NET EP FAIR 005 - PE0000013 “FAIR” - PNRR M4C2 Investment 1.3, financed by the European Union - NextGenerationEU.
Appendices
accompanying the paper
“Multi-Scale Node Embeddings for Graph Modeling and Generation”
by R. Milocco, F. Jansen and D. Garlaschelli
Appendix A Non-Negative Logistic PCA
$\mathcal{L}\left(\mathbf{B},\mathbf{C}|\mathbf{A}\right):=$ $\sum_{i<j}a_{ij}\ln\left(\sigma_{ij}\right)+\left(1-a_{ij}\right)\ln\left(1-\sigma_{ij}\right)$ (23)
$=$ $\sum_{i<j}\min(x_{ij},0)-\ln(1+e^{-\left|x_{ij}\right|})-(1-a_{ij})x_{ij}$ (24)
$\partial_{b_{ik}}\mathcal{L}=$ $\sum_{j(\neq i)}\left(a_{ij}-\sigma_{ij}\right)b_{jk}$ (25)
$\partial_{c_{ik}}\mathcal{L}=$ $-\sum_{j(\neq i)}\left(a_{ij}-\sigma_{ij}\right)c_{jk}$ (26)
$\partial_{b_{ik}}\mathcal{L}\stackrel{{\scriptstyle!}}{{=}}0$ (27)
$\partial_{c_{ik}}\mathcal{L}\stackrel{{\scriptstyle!}}{{=}}0$ (28)
@@ Appendix A Non-Negative Logistic PCA
A.1 Renormalizing the LPCA
The LPCA does not have a recipe to renormalize the parameters and produce an “up-scaled” version of it. Nevertheless, it is possible to model the multi-level structure $\left\{\mathbf{A}_{\ell}:\ell\geq 0\right\}$ either through the coarse-graining $\mathbf{P}_{cog}^{(\ell)}$ or by fitting it at every level. Both of them suffers from the lack of scale-invariance, as we show in the following.
Firstly, one may calculate $\mathbf{P}_{cog}^{(\ell)}$ from the RHS of the Equation 34, namely
$p^{cog,(\ell)}_{IJ}:=1-\prod_{i_{0}\in I,j_{0}\in J}\frac{1}{1+e^{\langle\vec{b}_{i_{0}},\vec{b}_{j_{0}}\rangle-\langle\vec{c}_{i_{0}},\vec{c}_{j_{0}}\rangle}}$ (29)
where $I:=\mathbf{\Omega}_{0\to\ell}(i_{0}),J:=\mathbf{\Omega}_{0\to\ell}(j_{0})$ are the block-nodes at level $\ell$.